OpenStudy (anonymous):

how much would $500 invested at 4% interest compounded continuously be worth after 7 years?

OpenStudy (anonymous):

$$$ = P(1+i)^n $$$ = current total P = principal i = interest n = length of time $$$ = 500 * (1 + 0.04)^7 $$$ = 657.97